OpenStudy (anonymous):

A wheel of mass 15kg and radius 1m is being pulled by a horizontal force F against a step 0.4m high. What initial force is jut sufficient to turn the wheel so that it will rise over the step?

OpenStudy (anonymous):

Do u have answer key with u

OpenStudy (anonymous):

i got 120N

OpenStudy (anonymous):

It's supposed to be 200N

OpenStudy (anonymous):

Hmm, I'm getting about 250N . . .

OpenStudy (anonymous):

Hmm

OpenStudy (anonymous):

Are we assuming that the force is being applied through the center of the wheel?

OpenStudy (anonymous):

Ah, i think I know what I did wrong. The length of the lever arm is changing as the wheel climbs the step...

OpenStudy (anonymous):

mm torques should be equal .. look at the step edge now the torques are due to Mg and F  length of A is 1-0.4 = 0.6 length of B is sqrt(1^2 - 0.6^2) = 0.8 so 0.6 * F = 0.8 * 150 F = 200